### v2

# Leopard[Uncaptioned image]: A Vision Language Model for Text-Rich Multi-Image Tasks

Abstract
Text-rich images, where text serves as the central visual element guiding the overall understanding, are prevalent in real-world applications, such as presentation slides, scanned documents, and webpage snapshots. Tasks involving multiple text-rich images are especially challenging, as they require not only understanding the content of individual images but reasoning about inter-relationships and logical flows across multiple visual inputs. Despite the importance of these scenarios, current multimodal large language models (MLLMs) struggle to handle such tasks due to two key challenges: (1) the scarcity of high-quality instruction tuning datasets for text-rich multi-image scenarios, and (2) the difficulty in balancing image resolution with visual feature sequence length. To address these challenges, we propose Leopard, a MLLM designed specifically for handling vision-language tasks involving multiple text-rich images. First, we curated about one million high-quality multimodal instruction-tuning data, tailored to text-rich, multi-image scenarios. Second, we developed an adaptive high-resolution multi-image encoding module to dynamically optimize the allocation of visual sequence length based on the original aspect ratios and resolutions of the input images. Experiments across a wide range of benchmarks demonstrate our model’s superior capabilities in text-rich, multi-image evaluations and competitive performance in general domain evaluations. Our code is available at https://github.com/Jill0001/Leopard.
1 Introduction
Figure 1: Left: A demonstration of a text-rich multi-image task. Models need to reason about the textual content across multiple images to answer the question correctly. Leopard  successfully generates the right answer while baselines fail. Right: Evaluation results of Leopard and three baselines. Our model surpasses its counterparts across text-rich multi-image benchmarks by a large margin, maintaining comparable performance on single and general evaluations.
Multimodal large language models (MLLMs) have revolutionized vision-language tasks, driving advancements in a variety of areas such as image captioning and object detection [67, 74, 72]. These improvements extend to applications involving text-rich images where text serves as the primary visual element guiding image comprehension, such as visual document understanding [51] and scene text recognition [59]. Traditional OCR-based pipelines in these text-rich visual scenarios are being replaced by end-to-end approaches that directly encode intertwined multimodal inputs [69, 75, 63], leading to improved efficiency and accuracy in handling text-rich images.
Despite these advancements, the majority of existing open-source MLLMs, like LLaVAR [75] and mPlug-DocOwl-1.5 [16], have primarily focused on optimizing performance for text-rich single-image tasks. This focus inherently limits their applicability in many real-world scenarios, where tasks often involve multiple inter-connected images. For instance, multi-page visual document understanding requires integrating information spread across different pages to capture the logical flow across the whole document [64, 25]. To understand presentation slides, grasping the overarching narrative requires understanding multiple slides with unique but interrelated content [62]. These vision-language tasks on multiple text-rich images require advanced capabilities that go beyond merely recognizing text and visuals within a single image; they involve understanding and reasoning about relationships and logical flows across multiple visual inputs. While some models – such as OpenFlamingo [3], VILA [38], Idefics2 [28] – have made strides toward supporting multi-image inputs, they mainly focus on scenarios with natural images but fall short in understanding sequences of text-rich images with interrelated textual and visual information. We plot the performance of representatives of the aforementioned models in Figure 1. Upon examining their training data and model architecture, we identified two primary limitations.
First, there is a scarcity of high-quality instruction tuning datasets on text-rich multi-image scenarios. Existing visual instruction tuning datasets for text-rich images are predominantly based on single-image inputs [22, 59, 48, 63], which limits the model ability to generalize and reason across multiple images. Second, in text-rich multi-image scenarios, there is a challenge of balancing image resolution and sequence length limitations. Many general-domain MLLMs adopt the low-resolution settings of pre-trained visual encoders [38, 21]. However, for text-rich images, such as scientific reports, recognizing text content becomes difficult at low resolutions. While some approaches overcome this in single-image settings by splitting the original image to preserve high-resolution details [40, 16], this approach is less effective when applied to multiple images, as it quickly exceeds model’s maximum sequence length. Moreover, compressing such long-sequence representations into shorter ones leads to significant information loss, thereby degrading model performance [3, 27]. Thus, a critical balance must be struck between maintaining sufficient visual detail and keeping sequence lengths manageable.
In this paper, we introduce a novel multimodal large language model, named Leopard [^1] Leopards have remarkable visual adaptations that allow them to track prey both from afar and up close, making them highly efficient hunters.. Leopard is specifically designed to handle complex text-rich, multi-image tasks. To train Leopard, we first curated about one million high-quality multimodal instruction-tuning data, tailored to the text-rich, multi-image scenarios. This dataset spans three key domains that are commonly encountered in real-world scenarios: (1) multi-page documents, (2) multi-charts and multi-tables, (3) webpage trajectories. These scenarios capture the increasing complexity and multimodal nature of modern digital information. In addition, to enable high-resolution encoding in multi-image inputs, we equipped Leopard with an adaptive high-resolution multi-image encoding module. Specifically, it dynamically optimizes the allocation of visual sequence length based on the original aspect ratios and resolutions of the input images. We then apply pixel shuffling to losslessly compress [8] long visual feature sequences into shorter ones. This approach allows the model to accommodate multiple high-resolution images without compromising detail or clarity.
We conducted experiments on 13 vision-language benchmark datasets, evaluating Leopard from multiple perspectives. Consistent improvements were observed when training Leopard with two distinct base model architectures: LLaVA and Idefics2. Our results demonstrate Leopard’s superior performance on 5 text-rich, multi-image benchmarks, outperforming the best open-source MLLM by an average of +9.61 points. Moreover, Leopard remains highly competitive in text-rich single-image tasks and general-domain vision-language benchmarks, achieving comparable results to state-of-the-art MLLMs without extensive fine-tuning. Further ablation studies confirm the effectiveness of our instruction-tuning dataset and the adaptive high-resolution encoding module. These findings highlight Leopard’s strong performance across various multimodal applications.
2 Related Work
Multimodal Large Language Models (MLLMs). Many approaches have been proposed for building MLLMs, leveraging different architectural designs. A widely adopted approach is the decoder-only architecture, exemplified by LLaVA [41], Emu2 [61], and Intern-VL [9]. These models typically incorporated a visual encoder to encode images, a vision-language connector to project visual features into the language feature space, and a language model that processes both visual and textual information jointly. Another line of work employed cross-attention architectures where encoded image features are integrated with textual tokens via cross-attention layers, as seen in Flamingo [1], OpenFlamingo [3] and CogVLM [66]. Such a design allows models to retain the benefits of a fully intact language model but introduces new parameters to manage the visual-textual interplay.
Text-rich MLLMs. Text-rich images are traditionally processed in pipelines [60, 18], where an OCR module first recognized text from the image, followed by processing through a language model. To improve efficiency and avoid error propagation, with the advent of MLLMs, end-to-end approaches become more popular recently. For instance, LLaVAR [75] utilized a dataset of 400K instances with OCR-enhanced text to outperform LLaVA on various text-rich VQA tasks. Subsequent models such as UReader [70], TextMonkey [44], and Mplug-DocOwl-1.5 [16] recognized the importance of high-resolution encoding for accurate text comprehension, so they adopted strategies that cropped single images into multiple sub-images to preserve the original resolution during visual encoding. However, these approaches are primarily trained on single-image data, and struggle to generalize effectively to multi-image scenarios. Furthermore, the straightforward partitioning technique encounters challenges with multi-image inputs, as the sequence length rapidly increases with the number of images.
Multi-image MLLMs. Efforts have been made in training MLLMs with multi-image inputs due to the prevalence of multi-image scenarios in real-world applications. Mantis [21] introduced a multi-image instruction tuning dataset on a variety of natural image scenarios. Besides, both VILA [38] and Idefics-2 [28] incorporated image-text interleaved data during their pre-training. LLaVA-Next-Interleave [33] further extended this by incorporating videos and multi-view 3D data into the training pipeline. However, these works primarily target natural images and general visual understanding, leaving a gap in handling text-rich, multi-image scenarios. Natural images typically follow a different distribution from text-rich images and often do not demand high-resolution processing. As a result, many existing multi-image MLLMs struggle to generalize to text-rich scenarios. Our work aims to address this gap by specifically focusing on multi-image settings where text-rich images are the primary input. Very recently (in 08/2024 and 09/2024), multi-image training for MLLMs has attracted intense attention from researchers. Several concurrent efforts have included multi-image interleaved data to train their models, such as LLaVA-OneVision 08/2024 [31], Idefics3 (08/2024, 26), NVLM (09/2024, 10), mPlug-DocOwl-2 (09/2024, 17), Molmo (09/2024, 11) and Qwen2-VL (09/2024, 65). This trending paradigm highlights the significant practical value of multi-image MLLMs by enhancing their ability to tackle a wide range of real-world applications. The incorporation of multi-image instruction tuning data is therefore of paramount importance.
Figure 2: The overall model pipeline. Given ① raw image inputs, ② we first compute the optimal allocation of sub-image numbers and splitting strategy for all images based on their resolution and aspect ratio. ③ The images undergo padding, resizing, and splitting operations. ④ Both sub-images and resized original images are then encoded into a sequence of visual features. These sequences subsequently undergo a pixel shuffle operation that concatenates every four features. ⑤ The visual features are projected into the language embedding space via a vision-language connector. Finally, the large language model then integrates these visual and language embeddings to generate responses.
3 Method
Leopard follows the typical design of decoder-only vision language models [41, 40, 33], including a visual encoder, a vision language connector, and a language model (LM), as shown in Figure 2 (④⑤). Specially, the input images are first passed through the visual encoder, which extracts high-level visual features and captures essential semantic information. These visual features are then projected into the language representation space via the vision-language connector. After this transformation, the visual tokens are interleaved with the textual tokens, resulting in a sequence of interleaved text-visual tokens. This interleaved sequence is then fed into the LM, which processes these inputs in a causal manner, leveraging the contextual dependencies between text and visual information to generate coherent outputs that align with both modalities.
3.1 Multi-image text-rich Instruction Turning
Dataset
To train Leopard, we construct a large instruction-tuning dataset named Leopard-instruct, comprising 925K instances, with 739K specifically designed for text-rich, multi-image scenarios. While we extensively surveyed existing open-source datasets, we only identified 154K usable text-rich, multi-image samples, which is far from sufficient for effective instruction tuning, as shown in prior MLLM studies [21, 28, 33]. To address this data scarcity, we developed several data collection pipelines to collect high-quality text-rich, multi-image data, resulting in additional 585K instances. Each instance consists of a set of images along with corresponding task instructions and responses. The dataset details are presented in Table 1, and a detailed breakdown of its composition can be found in Appendix A.1.
Documents and Slides are common sources of multi-image data that primarily contain text and require cross-page context integration to fully understand the information.
These data is collected in three ways. First, we include 69K public multi-page document and slide datasets [64, 25, 79, 62], covering a variety of document types such as scanned handwriting, printed documents, and digital PDFs. Second, we adapt two single-page document datasets, DocVQA [51] and ArxivQA [34], for multi-image settings. Following [21], we randomly merge 2 to 4 single-page instances by concatenating their respective images and Q-A pairs. Prompts like “in the second image” are added to direct the model’s focus to the appropriate image. These merged samples help the model learn how natural language references align with corresponding image features. Third, we collect raw slides from [57] and SlideShare [^2] https://www.slideshare.net, and use GPT-4o to generate Q-A pairs and reasoning steps. We show the prompt to GPT in Figure 5. Upon manually reviewing 100 instances annotated by GPT-4o, we found an accuracy rate over 90%, indicating high annotation quality.
Tables and Charts provide highly organized, structured quantitative information, often involving complex data patterns and relationships, requiring the integration of both visual and textual elements for accurate interpretation.
To address the lack of instruction tuning data involving multiple tables or charts, we use the following strategies. First, we include 21K open-source multi-chart and multi-table datasets [77, 54], originally stored in JSON or DataFrame formats. We programmatically render these tables as images, converting them into multimodal data. Details of rendering can be found in Appendix A.3 Second, We utilize the TableGPT [35] dataset and split each table into multiple sub-tables, then convert them into figures, thereby creating multi-modal, multi-table instruction data. Third, we apply the same merging strategy used for combining single-page documents to synthesize multi-image datasets. This approach integrates several single-chart datasets, including ChartGemma [49], ChartQA [48], DVQA [22], and FigureQA [23]. Besides, we generate new multi-chart data from social reports of the Pew Research Center [^3] https://www.pewresearch.org that feature multiple interrelated charts within the articles under the same topic. We download charts from the website and use GPT-4o to create 20K Q-A pairs that require multi-chart understanding.
Data Types # Instances Proportion
Total Samples 925K
Single-image 186K 20.10%
Multi-image 739K 79.89%
    *Public 154K 16.65%
    *New (Ours) 585K 63.24%
Rationales
    *Existing 214K 23.14%
    *New (Ours) 250K 27.02%
    *None 461K 49.84%
Domains
     Documents 192K 20.76%
     Slide Decks 16K 1.73%
     Tables 48K 5.19%
     Charts 353K 38.16%
     Webpages 55K 5.95%
     Others 261K 28.22%
Table 1: Data statistics of the Leopard-Instruct dataset.
Webpage Snapshots consist of sequential images representing web pages, providing visual context for user interactions and task flows. Understanding webpage is a critical skill for MLLMs to evolve into fully autonomous web agents [12, 14]. To collect relevant data, we format several web-related multimodal datasets into a Q-A structure as follows:
  1. 1.
    Web action prediction data: We include Mind2Web [12] and OmniACT [24], where we divide long web snapshots into multiple sub-figures, and plot bounding boxes based on the coordinates of web elements. Then GPT-4o is used to convert the original action data into a Q-A format, where the task is to identify the correct element to interact with.
  2. 2.
    Web-based classification data: We incorporate WebScreenshots [4], WebVision [36], and WebUI [68]. We utilize the web snapshots in these datasets and employ GPT-4o to generate Q-A pairs on webpage understanding, including chain-of-thought reasoning steps. The prompting details are provided in Figure 6.
Models Visual Encoder Resolution Backbone LLM Param. PT. IT.
Otter-9B [30] CLIP ViT-L $224^{2}$ LLaMA-7B 9B 30M 5.1M
Emu2-Chat [61] EVA-02-CLIP $448^{2}$ LLaMA-33B 37B - 160M
MM1-7B-Chat [52] CLIP ViT-H $378^{2}$ 7B - 1.5M
VILA1.5-8B [38] SigLIP $384^{2}$ LLaMA3-8B 8B 50M 1M
mPlug-DocOwl-1.5 [16] CLIP ViT-L $448^{2}$ (x9 crops) LLaMA-7B 8B 4M 1M
Idefics2-8B [28] SigLIP $980^{2}$ Mistral-7B 8B 350M 20M
LLaVA-NeXT-Inter [33] SigLIP AnyRes Qwen1.5-7B 7B 1.3M 1.2M
Mantis-LLaVA [21] SigLIP $384^{2}$ LLaMA3-8B 8B 0.5M 1M
Mantis-Idefics2 [21] SigLIP $980^{2}$ Mistral-7B 8B 350M 1M
Leopard-LLaVA (Ours) SigLIP Adapt HR. LLaMA3.1-8B 8B 0.5M 1.2M
Leopard-Idefics2 (Ours) SigLIP $980^{2}$ Mistral-7B 8B 350M 1.2M
Table 2: A detailed comparison of the model training details, including image resolution, vision encoder, backbone LLM, number of parameters (Param.), pre-training (PT.) data size, and instruction tuning (IT.) data size of baselines. AnyRes denotes the resolution selecting method proposed by [40] and Adapt HR. represents the proposed adaptive high-resolution multi-image encoding strategy.
Augmenting with Rationales. In contrast to single-image tasks, multi-image scenarios typically require MLLMs to integrate information across multiple images, making cross-image reasoning difficult to train when only the final answer is provided [78, 19]. To address this, we employ GPT-4o to generate chain-of-thought (CoT) rationales for inherently multi-image datasets (excluding those formed by merging single-image data) that lack CoT annotations. This results in 250K instances with GPT-annotated reasoning, with the prompt detailed in Figure 7.
Other Domains. We also include datasets from various other domains such as maps (MapQA, 6), infographics (InfographicVQA, 50), mathematical diagrams (MathV360K, 58), and abstractive diagrams (IconQA, 47). We also incorporate mixed-domain datasets for text-rich images, including LLaVAR [75], Monkey, 37, and mPlugDocReason [16]. We remove duplicate subsets from these mixed-domain datasets. Among these datasets, 64K samples consist of multi-image data, while the remaining are single-image samples. To preserve natural image understanding ability, we add 313K samples from ShareGPT4V [7], an instruction dataset for natural images.
3.2 Adaptive High-resolution Multi-image
Encoding
Image resolution significantly influences the visual perception and understanding capabilities of MLLMs, particularly when processing text-rich images. Low-resolution images often cause printed text to become blurred or unreadable, resulting in misinterpretations, perception errors, and visual hallucinations. The visual resolution of most existing MLLMs is determined by their pre-trained visual encoders, which are typically limited to low resolutions such as $224\times 224$ or $336\times 336$ pixels [39, 38, 21]. These low-resolution constraints can hinder MLLMs to accurately understand textual information embedded within images.
To overcome these limitations, a natural solution is dividing a high-resolution image into multiple smaller sub-images, each of which is independently processed by the model’s visual encoder [40, 13]. This partitioning allows for the extraction of more fine-grained visual details, making it possible to capture small or densely packed textual elements. However, a major drawback of this approach is that it significantly increases the length of visual feature sequence. When applied to scenarios involving multiple image inputs, the feature sequences are easily exceeding the model’s maximum sequence length limit. To address the issue, we follow the image-splitting idea and propose a novel adaptive high-resolution multi-image encoding strategy as follows.
[…]
$S_{i}=\left\lfloor\frac{h_{i}}{v}\right\rfloor\times\left\lfloor\frac{w_{i}}{v}\right\rfloor,$ (1)
where $v$ is the resolution of visual encoder (e.g., $v=364$ pixels). If the total number of patches satisfies $\sum_{i}S_{i}\leq M$, we proceed with these sub-image counts. Otherwise, we scale down these counts proportionally using a scaling factor $\alpha=\frac{M}{\sum_{i}S_{i}}$, resulting in adjusted sub-image counts:
$S^{\prime}_{i}=\left\lfloor\alpha S_{i}\right\rfloor.$ (2)
Image Partitioning: For each image, we perform a grid search over possible number of rows $r$ and columns $c$ (where $1\leq r,c\leq S^{\prime}_{i}$ and $r\times c\leq S^{\prime}_{i}$) to find the optimal cropping configuration that maximizes the effective resolution within the allocated sub-images [32]. This configuration results in the original image being padded and resized to a target resolution of $(h^{\prime}_{i}=r\times v,w^{\prime}_{i}=c\times v)$. We then divide the image into $r\times c$ sub-images of size $(v\times v)$. Additionally, the original image is directly resized to $(v\times v)$, which provides a global view of the visual content.
[…]
4 Experiment
Models Text-Rich Multi-Image Text-Rich Single-Image
MVQAD DUDE SlideVQA MCQA MH Multi Avg. VQAT VQAD VWB Avg.
Otter-9B 0.17 0.15 5.95 1.08 0.14 1.50 23.18 3.53 10.20 12.30
Emu2-Chat 17.58 13.79 0.60 2.40 0.72 7.02 66.60 5.44 18.17 30.07
MM1-7B-Chat - - - - - - 72.80 - - -
VILA-LLaMA3-8B 30.75 19.75 24.72 1.87 3.66 16.15 66.30 30.38 23.37 40.02
mPlug-DocOwl-1.5 35.85 16.94 4.54 0.26 0.86 11.69 68.60 82.20 29.80 60.20
Idefics2-8B 46.67 23.06 25.14 2.59 9.89 21.47 70.40 67.30 23.76 53.82
LLaVA-NeXT-Inter 39.92 24.04 23.46 14.34 3.55 21.06 62.76 75.70 21.36 53.27
Mantis-LLaVA 31.89 17.73 16.81 9.72 3.46 15.92 59.20 39.02 17.88 38.70
Mantis-Idefics2 51.61 27.74 24.02 12.97 5.48 24.36 63.50 54.03 22.47 46.67
Leopard-LLaVA 53.90 35.94 23.83 9.68 10.76 26.82 67.70 68.07 24.91 53.56
Leopard-Idefics2 66.06 40.74 34.93 18.03 10.09 33.97 80.40 74.79 25.60 60.26
Table 3: Experiment results of baseline models and Leopard on 8 benchmarks of text-rich images. We use abbreviated benchmark names due to space limits. MVQAD: Multi-page DocVQA, MCQA: MultiChartQA, MH: MultiHiertt, VQAT: TextVQA, VQAD: DocVQA, VWB: VisualWebBench. Following [64], for MVQAD, DUDE, and VQAD, we use average normalized levenshtein similarity [5] as the metric. For other benchmarks, accuracy is used as the metric, which measures if the predicted answer exactly matches any target answer.
4.1 Implementation Details
Model Architecture. We train our models on two base architectures: LLaVA [39] and Idefics2 [28]. For Leopard-LLaVA, we use SigLIP-SO-400M [73] with $364\times 364$ image resolutions as the visual encoder since it supports larger resolution than the commonly used $224\times 224$ resolution CLIP visual encoder [55]. Each image is encoded into a sequence of $26\times 26=676$ visual features under a patch size of $14$. With the visual feature pixel shuffling strategy, each image is further processed into a sequence of $169$ visual features. We limit the maximum number of images ($M$) in each sample to $50$, which produces up to $8,450$ visual features in total. Following llava [39], we adopt a two-layer MLPs as the visual-language connector. We use LLaMA-3.1 [53] as the LM.
For Leopard-Idefics2, we follow the architecture of Idefics2-8B which uses SigLIP-SO-400M as the visual encoder but increases its image resolution to $980\times 980$ to make the text legible. The features outputted by the visual encoder are compressed with a feature resampler into $64$ tokens per image. Idefics2-8B adopts the Mistral-7B [20] as the LM.
Training Details. When training Leopard-LLaVA, we first train the visual-language connector using LLaVA’s 558K multimodal pre-training dataset. Subsequently, we fine-tune the model (with both the connector and the LM unfrozen) using our Leopard-instruct data. As for Leopard-Idefics2, it is pre-trained on a dataset comprised of over 350M multimodal samples. Given the computational challenges of reproducing such extensive pre-training, and to ensure a fair comparison with baselines that utilize the pre-trained Idefics2 checkpoint, we directly adopt Idefics2’ visual feature resampler and fine-tune the model on the Leopard-instruct dataset.
We train both Leopard-LLaVA and Leopard-Idefics2 on 64 A100-40G GPUs with a global batch size of 128. We use the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$. Following [21], we use a learning rate of $1\times 10^{-5}$ for Leopard-LLaVA and $5\times 10^{-6}$ for Leopard-Idefics2 to protect its pretrian knowledge. We use a cosine learning rate scheduler with a linear learning rate warm-up for the first $3\%$ steps. All model variants are trained 1 epoch under the same hyperparameters. It takes around 120 GPU days to train Leopard under both settings.
4.2 Baseline Models
We compare Leopard against a range of existing open-source MLLMs that support multi-image inputs. These baseline models include Otter-9B [30], Emu2-Chat-34B [61], MM1-7B-Chat [52], Mantis [21], VILA [38], Idefics2-8B [28], and LLaVA-NeXT-Interleave [33].
Models MIRB MiBench MMMU MathVista SQAI Avg.
Otter-9B 20.74 43.72 30.89 22.00 60.44 35.55
Emu2-Chat 36.02 58.93 34.10 30.40 65.69 45.03
MM1-7B-Chat - - 37.00 35.90 72.60 -
VILA-LLaMA3-8B 40.87 53.70 36.90 35.40 79.90 49.35
mPlug-DocOwl-1.5 25.39 40.80 35.44 29.50 64.40 39.11
Idefics2-8B 33.02 46.39 42.90 45.00 89.04 51.27
LLaVA-NeXT-Inter 44.38 74.52 38.44 32.10 72.63 52.41
Mantis-LLaVA 40.76 59.96 40.10 34.40 74.90 50.02
Mantis-Idefics2 41.80 56.80 41.10 40.40 81.30 52.28
Leopard-LLaVA 42.00 60.80 43.00 45.50 85.57 55.37
Leopard-Idefics2 41.38 61.74 40.11 44.80 90.38 55.68
Table 4: Experimental results on general domain benchmarks. We abbreviate the $\mathrm{Image}$ split of ScienceQA as SQAI.
Models that only support a single image input are excluded from our comparisons, except for mPlug-DocOwl-1.5 [16], as it is primarily trained on visual document data and demonstrates strong capabilities on text-rich image tasks. Table 2 demonstrates a detailed comparison of the model training details of between baseline models and our proposed Leopard, which highlights their architecture, image resolution and training data differences.
4.3 Evaluating Benchmarks
We evaluated Leopard and baseline methods across three categories of vision-language tasks on (1) single text-rich image evaluation, (2) multiple text-rich images evaluation, and (3) general reasoning evaluation. Benchmarks for (1) include TextVQA [59], DocVQA [51], and VisualWebBench [43]. Benchmarks for (2) include Multi-page DocVQA [64], DUDE [25], SlideVQA [62], Multihiertt [77], and MultiChartQA [2], which cover a diverse range of typical multi-image tasks, such as document understanding and slide question answering. Benchmarks for (3) include MMMU [71], MathVista [45], ScienceQA [56], MIRB [76] and MiBench [42], which evaluate MLLMs from different perspectives, including world knowledge, mathematics, and scientific reasoning etc.
Ablation Settings Text-Rich Multi-Image Text-Rich Single General
MVQAD DUDE SlidesVQA Multi Avg. TextVQA DocVQA MMMU MathVista
($\star$) Our Best Setting (as in Table 3): LLaMA-3.1 + Adaptive + 1 1 1
Leopard-LLaVA 53.90 35.94 23.83 37.89 67.70 68.07 43.00 45.50
(1) Effect of Adaptive High-Resolution Encoding: LLaMA-3.1 + 1 1 1
- w/o Adaptive 40.44 26.16 20.93 29.17(8.7$\downarrow$) 60.18 44.69 41.00 42.40
(2) Effect of Backbone LLMs: LLaMA-3 + Adaptive + 1 1 1
- with LLaMA-3 .1 48.66 32.64 25.75 35.68(2.2$\downarrow$) 67.08 54.92 41.22 42.10
(3) Effect of Data Domains: LLaMA-3.1 + Adaptive
- with chart web 43.79 29.50 23.10 32.13(5.7$\downarrow$) 66.78 56.60 40.67 44.80
- with doc web 54.33 35.65 18.73 36.23(1.7$\downarrow$) 66.86 50.78 41.89 39.60
- with doc chart 54.62 35.70 20.79 37.02(0.9$\downarrow$) 67.40 67.82 41.78 44.00
Table 5: Ablation studies on Leopard-LLaVA from four different perspectives: (1) evaluating the impact of Adaptive High-Resolution Encoding, (2) pre-training LLaVA by initializing with checkpoints from either LLaMA-3 or LLaMA-3.1, and (3) examining the impact of using different data domains for instruction tuning, including doc, chart, and web.
4.4 Main Experimental Results
Question 1: How does Leopard compare to state-of-the-art MLLMs on vision-language tasks?
Leopard achieves outstanding performance on text-rich, multi-image benchmarks, as shown in Table 3. Notably, both Leopard-LLaVA and Leopard-Idefics2 significantly outperform all baselines. Leopard-Idefics2 becomes the strongest open-source MLLM in this area, achieving an average improvement of 9.61 points over the previous best performance.
In single-image text-rich scenarios, Leopard outperforms several recent strong models, including VILA and LLaVA-NeXT. Leopard even achieves slightly higher average scores than the state-of-the-art mPlug model, despite mPlug being trained on 4M single-image data while Leopard is tuned on <200K. This demonstrates that training on multi-image data from Leopard-instruct also benefits model performance on single-image tasks.
In addition, we evaluate Leopard on general-domain benchmarks which contain both multi-image and single-image instances. As shown in Table 4, Leopard outperforms other open-source MLLMs on these benchmarks. Remarkably, Leopard surpasses Mantis, its counterpart multi-image model trained on the same foundational architecture and a comparable volume of data. This performance demonstrates the high quality and diversity of the Leopard-instruct dataset, which effectively preserves our model’s general image understanding capabilities.
Question 2: Is the one-million text-rich multi-image dataset effective for instruction tuning?
Mantis-Idefics2 is trained on a combination of natural multi-image data and text-rich single-image data. However, Leopard-Idefics2 outperforms Mantis-Idefics2 by 12.8 points on text-rich multi-image benchmarks. This disparity indicates that developing strong multi-image text-rich capabilities through cross-domain transfer, such as with Mantis data, presents significant challenges. This finding underscores the importance of optimizing Leopard using high-quality, diverse, and well-curated multi-image text-rich datasets that are specifically tailored for complex multi-image scenarios.
Furthermore, Leopard-Idefics2 surpasses its base model, Idefics2, by 6.4 points across three single-image text-rich benchmarks, though Idefics2 is trained on over 20M instruction data that includes text-rich tasks like DocVQA and TextVQA. This highlights that the Leopard-instruct provides unique advantages to MLLMs that are not adequately addressed by existing datasets.
Question 3: Does Adaptive high-resolution multi-image encoding improve MLLM performance?
To assess the effectiveness of the proposed adaptive high-resolution multi-image encoding, we compared Leopard with a variant that excludes this feature (i.e., w/o Adaptive in Table 5). We notice a significant performance decline across all text-rich benchmarks, particularly on document-related benchmarks like DocVQA (-23.4), Multi-page DocVQA (-13.5), and DUDE (-9.8). This observation supports our hypothesis that high-resolution image encoding is especially beneficial for text-rich images, particularly with dense text content such as document pages.
4.5 More Analysis
Question 4: How does data from different domains contribute to instruction tuning?
Leopard-instruct mainly cover three main domains, i.e. documents & slides ( doc), tables & charts ( chart), and websites ( web). To assess the impact of data from different domains, we conduct ablation studies on three variants of Leopard, with the results presented in Table 5 Removing any part of the training data results in performance degradation. The most significant drop occurs when we exclude document data while removing web data leads to a slight decrease. However, the mixed-domain datasets, such as LLaVAR and mPlugDocReason, also contain data in these domains which are challenging to isolate and ablate. This may contribute to the relatively preserved performance even after the ablation of certain data sources.
Figure 3: Impact of the sub-image budget $M$ on the resulting model across four benchmarks. w/o indicates no partitioning into sub-images.
Question 5: What is the influence of different image budgets in adaptive multi-image encoding?
In our adaptive multi-image encoding module, we define a budget $M$ for the maximum number of sub-images that the model can process. To evaluate the impact of such image partitioning, we train Leopard using different values of $M$: 25, 50, 75, as well as a baseline setting where no image partitioning is applied and the number of sub-images equals the number of original images. According to the results plotted in Figure 3, model performance peaks or plateaus when $M$ is set around 50. Thus, we adopt 50 as the default value for training Leopard. These results show that increasing image numbers does not consistently improve performance, as input sequences can become excessively long and even exceed the model’s sequence length limit.
Question 6: How does the backbone language model affect the performance?
To ensure a fair comparison with multi-image competitor models, Mantis-LLaVA and VILA1.5, we also evaluate a variant of Leopard using LLaMA-3 instead of LLaMA-3.1, aligning its backbone language model architecture with these two baselines. According to Table 5, this substitution results in only a slight drop in average performance on text-rich multi-image tasks (2.2$\downarrow$). Nevertheless, comparing with results in Table 3, Leopard-LLaMA-3 still substantially outperforms both baselines in all tasks, such as Multi-page DocVQA (+16.8 over Mantis and +17.9 over VILA) and DUDE (+14.9 over Mantis and +12.9 over VILA). These results indicate that Leopard’s superior performance is not simply a result of the upgraded backbone large language models.
5 Conclusion
In this paper, we introduce Leopard, a novel MLLM specifically designed for text-rich, multi-image tasks. Leopard is equipped with two key innovations: (1) Leopard-instruct, a large-scale instruction-tuning dataset that encompasses a wide range of text-rich, multi-image instructions, and (2) an adaptive image encoding module capable of processing multiple high-resolution images efficiently. Our experimental results across diverse benchmarks highlight Leopard’s superior performance compared to existing open-source MLLMs, particularly in text-rich multi-image scenarios. Further analysis and ablation studies underscore the effectiveness of both the collected dataset and adaptive encoding strategy, solidifying Leopard’s contribution to multimodal research.
Appendix A Appendix
A.1 Leopard-Instruct
To train Leopard, we created a large instruction-tuning dataset, Leopard-instruct, with 925K instances, including 739K designed for text-rich, multi-image scenarios. Despite surveying existing datasets, we found only 154K suitable text-rich, multi-image samples – insufficient for effective instruction tuning, which is far from sufficient for effective instruction tuning, as shown in prior MLLM studies [21, 28, 33]. To overcome this limitation, we developed several data collection pipelines to collect high-quality text-rich, multi-image data, resulting in additional 585K instances.
Table 6 provides a detailed breakdown of the composition of the Leopard-instruct dataset. This table includes the name, domain, and sample size of sub-datasets. Additionally, it specifies how we construct multi-image samples, the number of images per sample, and the presence of rationales.
Dataset Domain Multi-image Images Rationales #Samples (K)
ArxivQA [34] Doc Reformed 1-3 Existing 81
DUDE [25] Doc Public 1-50 Augmented 23
MP-DocVQA [64] Doc Public 1-20 Augmented 36
DocVQA [51] Doc No 1 None 39
TAT-DQA [79] Doc Reformed 2-5 Augmented 13
SlidesGeneration [57] Slides Repurposed 1-20 Augmented 3
SlidesVQA [62] Slides Public 20 Augmented 10
Slideshare Slides Collected 2-8 Augmented 3
Multihiertt [77] Table Public 3-7 Existing/Augmented 15
MultiTabQA [54] Table Public 1-2 Augmented 6
TableGPT [35] Table Split 2 Existing 4
TabMWP [46] Table No 1 Existing 23
ChartGemma [49] Chart Reformed 1-4 Existing 65
DVQA [22] Chart Reformed 1-3 None 200
FigureQA [23] Chart Reformed 1-2 None 36
ChartQA [48] Chart Reformed 2 Augmented 32
Pew_MultiChart Chart Collected 2 Augmented 20
Mind2Web [12] Web Split 1-5 None 7
WebsiteScreenshots [4] Web No 1 Augmented 2
Omniact [24] Web No 1 None 1
RICO [15] Web Reformed 1-4 None 25
WebVision [36] Web No 1 Existing 1
WebUI [68] Web No 1 None 19
LLaVAR [75] Mix No 1 Existing 15
MathV360k [58] Mix No 1 None 38
Monkey [37] Mix Reformed 1-3 None 92
MPlugDocReason [16] Mix No 1 Existing 25
IconQA [47] Other Public 1-6 Augmented 64
InfographicVQA [50] Other No 1 Augmented 23
MapQA [6] Other Reformed 1-2 None 4
Total - - - - 925
Table 6: Details of the constructed Leopard-instruct dataset. Images denotes the image number of one sample in each dataset.
We draw a chart to illustrate the data composition of Leopard-instruct dataset 4.
Figure 4: An illustration of the proportion of sub-datasets and domains in the proposed dataset.
A.2 Prompts
We specify the prompt used during the data construction process as follows:
Figure 5: The prompt used for generating Q-A pairs with rationales for slide decks data.
Figure 6: The prompt used for generating Q-A pairs with rationales for webpage data.
Figure 7: We use this prompt for the generation of chain-of-thought rationales given original question, answer, and images.
A.3 Details of Table Rendering
To convert the textual table dataset into a multimodal dataset, the JSON or DataFrame format data is transformed into tabular images using Python. We utilize three Python packages, i.e. dataframe_image [^5] https://github.com/dexplo/dataframe_image., pandas [^6] https://pandas.pydata.org/., and matplotlib [^7] https://matplotlib.org/. with various styling to enhance the diversity of the rendered images. To ensure the clarity and legibility of the plotted images, the original data is filtered by excluding any tables that contain more than 20 rows. This threshold was set to maintain the recognizability of the resulting images.
A.4 Qualitative Results
We show two examples to give an illustrative demonstration of the model’s performance. As can be seen from Figure 8, Leopard can not only capture detailed data in multiple tables precisely but also perform cross-table calculations, therefore it can answer the complex question correctly. Another example is demonstrated in Figure 9. Leopard can accurately perceive the prominent information under a high-resolution four-page document, demonstration effective text-rich abilities under multi-image scenarios.
Figure 8: An example of multi-table reasoning of Leopard.
Figure 9: An example of multi-page document question answering of Leopard.